if w=64*63*62*61*......*3*2*1, which of the following choices is not a factor of w? a)69, b)68, c)67, d)66
Thank you

Answer by Edwin McCravy(20062)

if w=64*63*62*61*......*3*2*1, which of the following choices is not a factor of w? a)69, b)68, c)67, d)66
69 = 3*23
68 = 4*17
66 = 6*11

Those are all factor since 3,23,4,17,6 and 11 are all factors.
But 67 is prime and larger than 64 so it is the only one which 
is not a factor.

Edwin